# Reward Poisoning Attack Against Offline Reinforcement Learning

### v1

@@ Abstract
We study the problem of reward poisoning attacks against general offline reinforcement learning with deep neural networks for function approximation. We consider a black-box threat model where the attacker is completely oblivious to the learning algorithm and its budget is limited by constraining both the amount of corruption at each data point, and the total perturbation. We propose an attack strategy called ‘policy contrast attack’. The high-level idea is to make some low-performing policies appear as high-performing while making high-performing policies appear as low-performing. To the best of our knowledge, we propose the first black-box reward poisoning attack in the general offline RL setting. We provide theoretical insights on the attack design and empirically show that our attack is efficient against current state-of-the-art offline RL algorithms in different kinds of learning datasets.
Machine Learning, ICML
@@ 1 Introduction
Reinforcement learning is a popular learning paradigm for studying sequential decision-making problems (Sutton & Barto, 2018; Henderson et al., 2018), which are the essence of a majority of real-world applications. Specifically, reinforcement learning in the offline learning setting is more practical in many scenarios (Levine et al., 2020; Agarwal et al., 2020). Here, the training data is collected in an offline manner. For example, in the medical treatment problem, the offline data can be a recording of how a human doctor treats a patient. In contrast, collecting online data for the medical treatment problem would require the learner to treat patients during training, which is both dangerous and infeasible.
Despite theoretical understanding and empirical success in simulated environments, offline reinforcement learning faces a threat in the real world. In many applications, the offline data is usually based on human feedback (Zheng et al., 2018; Kiran et al., 2021). An adversary can poison the reward signal in the training dataset. When the data is provided by a third party, multiple adversaries can collude to corrupt different parts of the data to achieve the same goal of misleading the learning process. Current state-of-the-art offline reinforcement learning algorithms (Kumar et al., 2020; Fujimoto & Gu, 2021; Kidambi et al., 2020; Wu et al., 2019; Cheng et al., 2022) only consider ideal environments without such threats, which motivates us to investigate the vulnerabilities of these algorithms to such attacks. We focus on developing practical reward-poisoning attacks where the learning algorithm is no longer able to learn a high-performing policy under the attack. In general, a practical attacker should be black-box, that is, it should be oblivious to the learning algorithm and the training process of the learner. The corruption it applies to the reward signal in the dataset should also be limited. We formally introduce our criteria for an attack to be practical in Section 2.
Challenges: We find the following challenges for developing a practical reward poisoning attack:
@@
    Fixed training dataset: In the online learning setting, an attacker can mislead the learner to collect data that only covers states and actions with low long-term rewards so that the learner never observes how a high-performing policy would work making it fail to learn a high-performing policy (Xu et al., 2022). In contrast, the training dataset is fixed in the offline learning setting. The learner can always observe how an expert would behave in the dataset regardless of the attack. The attacker needs to convince the agent through reward corruption that certain expert behavior is bad even if it frequently occurs in the dataset with high true rewards that indicate its high performance.
@@
    No access to training process: A majority of online learning attacks need to observe the behavior of the learning algorithm during the training time (Zhang et al., 2020; Rakhsha et al., 2020), which is practical in that case because the learner needs to interact with the online environment. From such observations, the attacker can infer the current status of the learning algorithm, for example, the Q function it learns. However, in the offline setting, there is no interaction between the learner and the environment. The attacker has to infer the status of the learning agent through the training data only.
@@
    Limited learnable policies: To make offline learning efficient, a learning algorithm needs to be pessimistic (Levine et al., 2020). That is, if the behavior of a policy is barely covered by the dataset, then the algorithms should not output such a policy as it is unable to properly evaluate the policy. As a result, the policies that are not covered by the dataset can be fundamentally impossible to be learned by the learning agent. If the attacker wants to mislead the agent into certain bad behavior, it has to identify the set of policies that can be learned by the learner and can only choose bad behavior from that set.
  4. 4.
@@
    Black box attack: As discussed previously, a practical attacker should be oblivious to the learning agent. Without knowing the details of the offline learning algorithm, it would be impossible to predict the exact influence of a reward perturbation on the learning result of the algorithm, which is required by a majority of poisoning attack strategies against RL.
@@
Contributions: To address the challenges above, we adopt a basic assumption about efficient offline learning algorithms. Under the assumption, we adopt an attack framework where the attacker can approximate the learning outcome of an efficient learning algorithm while being oblivious to the learning algorithm or the training process. We derive theoretical insights on developing attacks that the attacker should make high-performing policies appear low-performing to the agent and vice versa. Based on the insights, we propose our main attack method: policy contrast attack, that can be used to assess the practical robustness of offline RL algorithms. To the best of our knowledge, this is the first black-box reward poisoning attack in the general offline RL setting. The core idea of our attack is to make the learner believe that the high-performing policies supported by the dataset are sub-optimal while some low-performing policies are optimal. We empirically test the efficiency of our attack on various standard datasets from the D4RL benchmark (Fu et al., 2020) learned by different state-of-the-art offline RL algorithms (Kumar et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2021). We find that our attack is efficient in a majority of cases with a limited attack budget. We also show that our attack is not sensitive to the choice of hyper-parameters. We hope that our work can inspire the development of more robust offline RL algorithms.
@@ 2 Preliminaries
Reward poisoning attack against offline RL: We consider a reward poisoning attack model where a malicious adversary can poison the reward signals in the offline dataset to mislead the learning agent. Formally, for the $i^{\text{th}}$ observation tuple $(s_{i},a_{i},s^{\prime}_{i},r_{i})$, the attacker can inject a perturbation $\Delta_{i}$ to the reward signal, then the corrupted observation tuple becomes $(s_{i},a_{i},s^{\prime}_{i},r_{i}+\Delta_{i})$. We denote $\mathbf{\Delta}={\Delta_{i},\ldots,\Delta_{N}}$ to be the corruption strategy of the attack.For a practical threat model, the attacker should have limited abilities. Formally, we consider the following constraints on the attacker.
@@
    No knowledge of the learning algorithm: There are a variety of efficient offline RL algorithms with very different learning strategies. Unless the learning agent announces its learning algorithm to the public, it would be impractical for the attacker to assume which learning algorithm the agent would choose. Therefore, the attacker should not have any detailed knowledge of the learning algorithm except for some basic assumptions that the algorithm is efficient when there is no attack.
@@
    Limited Budget: To make the attack stealthy, the attacker should make the corruption as small as possible. First, the attacker should limit the maximal perturbation on each reward signal $||\mathbf{\Delta}||_{\infty}=\max|\Delta^{i}|$, as an extreme value of reward would make the data point very suspicious. Second, the attacker should also limit the total amount of perturbation $||\mathbf{\Delta}||_{1}=\sum_{i=0}^{N}|\Delta^{i}|$. This is because poisoning the dataset would involve accessing the dataset collection process, and it will be suspicious if the attacker frequently accesses the process.
The general goal of the attacker is to make the learning agent learn a low-performing policy from the poisoned dataset. Formally, we adopt the following criteria for measuring the efficiency of an attack. Denote $B=||\mathbf{\Delta}||_{\infty}$ and $C=||\mathbf{\Delta}||_{1}$ to be the attack budgets required by the attack. Let $V=\mathbb{E}[J(\hat{\pi})]$ where $\hat{\pi}$ is the policy learned by the learning algorithm under the attack $\mathbf{\Delta}$, and the expectation is over the randomness in the learning process. An attack is more efficient if, with the same value of $B$ and $C$, it has a lower value of $V$ compared to other attacks.
3 Methods
3.1 Adversarial Reward Engineering
Here, we introduce the framework we adopt to solve the above attack problem. Given that the attacker is unaware of the learning agent, it is necessary to establish certain general assumptions that apply to the algorithms employed by the agent. Considering that a learning agent must ensure the efficiency of its learning algorithm when working with an uncorrupted dataset, it becomes practical for the attacker to make certain basic assumptions regarding the learning algorithm’s efficiency. One common characteristic of current offline RL algorithms is referred to as ’pessimism’ (Levine et al., 2020). In offline RL, a major challenge is known as ’distribution mismatch’. It says that if the distribution of trajectories generated by a policy is rarely covered by the dataset, the estimation of the policy’s performance based on that dataset will be inaccurate. Hence, the idea of ’pessimism’ aims to exclude policies that are inadequately covered by the dataset. To achieve this, state-of-the-art learning algorithms utilize diverse approaches and offer various theoretical guarantees under different assumptions (Kidambi et al., 2020; Fu et al., 2020; Fujimoto & Gu, 2021; Cheng et al., 2022; Wu et al., 2019). Consequently, exactly characterizing the learning efficiency of these algorithms with a universal mathematical formulation is difficult. Therefore, we adopt the following simple and fundamental assumption to assess the general efficiency of an offline RL algorithm
Assumption 3.1 ($\delta$-Efficient offline learning algorithm).
@@ 3.1 Adversarial Reward Engineering
Given an offline RL dataset $\mathcal{D}$ collected from an environment $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ with state-action distribution $\mu$. There exists a class of policies $\Pi_{\mu}$ that is supported by the distribution $\mu$. Let $\pi_{0}$ be the policy learned by the learning algorithm. For an efficient offline RL algorithm, there exists a small value $\delta>0$ such that $J_{\mathcal{R}}(\pi_{0})\geq\max_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)-\delta$. That is, the algorithm will learn the policy with nearly the highest performance on $\mathcal{M}$ among the supported policies.
For the rest of the section, we assume that the learning agents always use a $\delta$-efficient learning algorithm. Based on the assumption, similar to the previous work in the online learning setting (Xu et al., 2022), we can address the attack problem by solving a reward function engineering problem. We call the framework ‘adversarial reward engineering’, which is formally defined as below.
@@
Under the adversarial reward engineering framework, the attacker constructs an adversarial reward function $\widehat{\mathcal{R}}$. The corresponding corruption on the $i^{\text{th}}$ data $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$ satisfies $\Delta_{i}=\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}$.
In the adversarial reward engineering framework, it is equivalent to say the corrupted dataset is collected from the adversarial environment ^$\widehat{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\widehat{\mathcal{R}})$ with the same state-action distribution $\mu$ as the original dataset. Under Assumption 3.1, Theorem 3.3 generally characterizes the efficiency of an attack under the adversarial reward engineering framework.

### v2

# Universal Black-Box Reward Poisoning Attack against Offline Reinforcement Learning

@@ Abstract
We study the problem of universal black-boxed reward poisoning attacks against general offline reinforcement learning with deep neural networks. We consider a black-box threat model where the attacker is entirely oblivious to the learning algorithm, and its budget is limited by constraining the amount of corruption at each data point and the total perturbation. We require the attack to be universally efficient against any efficient algorithms that might be used by the agent. We propose an attack strategy called the ‘policy contrast attack.’ The idea is to find low- and high-performing policies covered by the dataset and make them appear to be high- and low-performing to the agent, respectively. To the best of our knowledge, we propose the first universal black-box reward poisoning attack in the general offline RL setting. We provide theoretical insights on the attack design and empirically show that our attack is efficient against current state-of-the-art offline RL algorithms in different learning datasets.
@@ 1 Introduction
Reinforcement learning (RL) in the offline setting (Levine et al., 2020; Agarwal et al., 2020) has become a promising framework for realizing reinforcement learning applications in practice. Offline RL avoids the necessity of potentially expensive online data collection and can work with offline data directly. Taking medical treatment problems as an example, collecting online data would require testing on humans, which is dangerous and infeasible. In contrast, abundant recordings of how a human doctor treats a patient are available for offline training.
A number of effective algorithms have been proposed for the offline RL problem and achieve promising empirical performance on simulated environments. (Bhardwaj et al., 2024; Kumar et al., 2020; Fujimoto & Gu, 2021; Kidambi et al., 2020; Wu et al., 2019; Cheng et al., 2022). However, a practical threat arises in the real world that is usually overlooked in the simulations. In many applications, the offline data is usually based on human feedback (Zheng et al., 2018; Kiran et al., 2021), making it possible for an adversary to poison the reward signal in the training dataset. For example, the learning agent may ask a third-party agent to provide offline data. If the third-party agent is dishonest, it may provide human data with subtle malicious changes that are hard to detect but can potentially cause unwanted results for the learning agent. The threat of poisoning attacks even draws attention in recent reinforcement learning from human feedback (RLHF) Baumgärtner et al. (2024); Wang et al. (2023); Wu et al. (2024), and people are concerned that such attacks may greatly influence the performance of current RLHF approaches. The state-of-the-art offline reinforcement learning algorithms mentioned above only consider ideal environments without such threats. A few works Li et al. (2024); Ye et al. (2023) investigate the settings with the poisoning attack. However, they only consider the attacker adopting some random attack strategies, which may not be enough to evaluate the robustness of an algorithm properly. Therefore, it is important to investigate the vulnerabilities of current algorithms to the poisoning attacks with more efficient attack strategies.
In Fig 1, we show the framework of the attack. Specifically, we focus on a universal black-box reward-poisoning attack with a limited budget such that under the attack, any originally efficient learning algorithm can no longer learn a high-performing policy. Such an attack setting is more realistic due to the following reasons:
@@
    Universal and Black box: First, it is not in the interest of the learning agent to reveal its algorithm to the attacker. Second, offline RL algorithms are very different in learning strategies and frameworks. Therefore, it is hard for the attacker to assume which kind of learning algorithm the agent will use. The attacker has to work without any knowledge of the algorithm used by the agent and expect the agent to choose any efficient algorithm arbitrarily.
@@
    Reward poisoning: In RL environments, the reward signals are usually provided by the environment, such as human users, while the state signals are determined by the agent. For example, in reinforcement learning from human feedback (RLHF), the state signals are the prompts the learner gives, while the reward signals are feedback from humans. Therefore, it is more feasible for the attacker to poison rewards.
@@
    Limited budget: Too much perturbation on the reward signals makes it easier for the learning agent to detect. So, it is more practical for the attacker to work with limited budgets and perturb the training process as little as possible.
We formally introduce our criteria for an attack to be practical in Section 3.
Challenges: We find the following challenges for developing a practical reward poisoning attack.
  1. 1.
    Fixed training dataset: In the online learning setting, an attacker can mislead the learner to collect data that only covers states and actions with low long-term rewards so that the learner never observes how a high-performing policy would work, making it fail to learn a high-performing policy (Xu et al., 2022). In contrast, the training dataset is fixed in the offline learning setting. This results in the famous ‘pessimistic’ learning (Levine et al., 2020) in offline RL, and the learning agent is inclined to stick to the behaviors covered by the dataset. The learner can always observe how an expert would behave (the action) in the dataset regardless of the attack. This contributes to the fundamental difficulty of misleading the learning agent.
  2. 2.
@@
    Universal and Black box attack: A practical attacker should be oblivious to the learning agent and universal efficiency against any algorithms it might use. First, it is challenging to list all possible efficient learning algorithms that an agent might use. Second, without knowing the details of an offline learning algorithm, it would be challenging to predict the exact learning result of the algorithm under the attack.
@@
Contributions: To address the challenges above, our insight is to characterize general efficient offline RL algorithms with the pessimistic learning feature. For a pessimistic algorithm, it will find nearly the best policy well covered by the dataset. Based on the insight, we construct an attack framework called the ‘adversarial reward engineering framework’ such that one can evaluate the efficiency of an attack against universal efficient algorithms in the framework. To find an efficient attack under the framework, the key idea of our attack is to find the high-performing policies covered by the dataset and make them appear low-performing to the agent and vice versa. In this case, the algorithm will believe that some poor-performing policies are near optimal. We build our main attack method, ‘policy contrast attack,’ based on the idea and theoretically analyze the insight behind it based on general assumptions. To the best of our knowledge, this is the first universal black-box reward poisoning attack in the general offline RL setting. We empirically test the efficiency of our attack on various standard datasets from the D4RL benchmark (Fu et al., 2020) learned by different state-of-the-art offline RL algorithms (Kumar et al., 2020; Fujimoto & Gu, 2021; Kostrikov et al., 2021). Our attack is efficient in most cases with a limited attack budget. We also show that our attack is not sensitive to the choice of hyper-parameters. One can use our attack to assess the practical robustness of offline RL algorithms. We hope that our work can inspire the development of more robust offline RL algorithms.
Figure 1: Reward poisoning attack framework.
2 Related Work
Reward poisoning attack in online RL: Most studies on reward poisoning attacks against RL focus on the online learning setting. Zhang et al. (2020a); Huang et al. (2017) investigate observation perturbation attack where the attacker modifies the state signals to adversarial examples during training. Rakhsha et al. (2020); Zhang et al. (2020b) theoretically investigate the attack problem in the online RL setting with tabular MDPs. They show that given a specific learning algorithm or an algorithm with certain learning guarantees, attack strategies exist such that the learning algorithm will learn a low-performing policy even if the attack only slightly perturbed the training process. However, many of these are white-box attacks, and it remains unknown how to scale these techniques to general RL environments that require function approximation with deep neural networks. In the more complicated general RL setting, Sun et al. (2020) empirically shows that current state-of-the-art DRL algorithms are vulnerable under the poisoning attack when the attacker is aware of the learning algorithm. Xu et al. (2022); Xu & Singh (2023) further shows that even in the black-box setting where the attacker is oblivious to both the environment and the learning algorithm, the data poisoning attack can achieve targeted and untargeted attack goals with limited budgets.
Reward poisoning attack and defense in offline RL: There is limited work on the poisoning attack in offline RL. Ma et al. (2019) theoretically studies the problem for tabular MDPs that cannot scale to the more complicated general RL setting considered in our work. Furthermore, their attack assumes that the learning agent can learn the globally optimal policy, which is not appropriate and practical for recent offline RL algorithms. Li et al. (2024) theoretically and empirically shows that the learning efficiency of an efficient offline RL algorithm is related to the dataset quality regardless of the reward signals, which is closely related to our assumption in Section 4. To defend against the attack, Zhang et al. (2022); Ye et al. (2023) design algorithms and theoretically prove their robustness. They also empirically test their algorithm against some specific random attack strategies. We empirically show that our attack is much more efficient than such random attacks for different learning scenarios. No attack exists for general offline RL except for the random attacks. Still, we also adapt a state-of-the-art online RL attack from Xu et al. (2022) to our setting as an additional baseline.
@@ 3 Preliminaries
Reward poisoning attack against offline RL: We consider a reward poisoning attack model where a malicious adversary can poison the reward signals in the offline dataset to mislead the learning agent. Formally, for the $i^{\text{th}}$ observation tuple $(s_{i},a_{i},s^{\prime}_{i},r_{i})$, the attacker can inject a perturbation $\Delta_{i}$ to the reward signal, then the corrupted observation tuple becomes $(s_{i},a_{i},s^{\prime}_{i},r_{i}+\Delta_{i})$. We denote $\mathbf{\Delta}={\Delta_{1},\ldots,\Delta_{N}}$ to be the corruption strategy of the attack.For a practical threat model, the attacker should have limited abilities. Formally, we consider the following constraints on the attacker.
@@
    No knowledge of the learning algorithm: There are a variety of efficient offline RL algorithms with very different learning strategies. Unless the learning agent announces its learning algorithm to the public, it would be impractical for the attacker to assume which learning algorithm the agent would choose. Therefore, the attacker should not have any detailed knowledge of the learning algorithm.
@@
    Limited Budget: To make the attack stealthy, the attacker should make the corruption as small as possible. First, the attacker should limit the maximal perturbation on each reward signal $||\mathbf{\Delta}||_{\infty}=\max|\Delta^{i}|$, as an extreme value of reward would make the data point very suspicious. Second, the attacker should limit the total amount of perturbation $||\mathbf{\Delta}||_{1}=\sum_{i=0}^{N}|\Delta^{i}|$. The reason is that too many slightly problematic data points can also make the agent realize that the dataset is unreliable.
Given a fixed budget, the general goal of the attacker is to make the learning agent learn a low-performing policy. We will formally formulate the attacker’s goal as an optimization problem in the next section.
4 Policy Contrast Attack
In this section, we formally define and formulate the universal black-box offline RL attack as an optimization problem. We consider the attacks that are universally effective against any data-driven learning algorithms of high learning efficiency. We highlight the challenges in finding the efficiency of an arbitrary attack and focus on a framework called the ‘adversarial reward engineering framework’ where the behavior of any efficient algorithm becomes predictable under any instance of the framework. Finally, we present our attack algorithm, ‘policy contrast attack,’ which is an efficient instance under the framework.
4.1 Universal Attack on General Offline Reinforcement Learning as An Optimization Problem
Here, we formally formulate the universal black-box offline RL attack as an optimization problem. Let $\mathcal{L}$ be a class of offline RL algorithms that a learning agent might use. Let $\pi^{\text{Alg}}(\mathcal{D}\oplus\mathbf{\Delta})$ be a random variable representing the policy learned by a learning algorithm Alg training on the dataset $\mathcal{D}$ with corruption $\mathbf{\Delta}$. By applying a limited amount of perturbation, the attacker should make the agent learn a low-performing policy no matter which algorithm it chooses from $\mathcal{L}$.
$\min_{\Delta_{1:N}}V$ (1)
[…]
$\|\mathbf{\Delta}\|_{0}\leq B,\|\mathbf{\Delta}\|_{1}\leq C$
The problem is not complete yet as we have not specified $\mathcal{L}$, the class of algorithms an agent might use. The problem becomes trivial if $\mathcal{L}$ includes all possible policies. In this case, the agent can choose an algorithm that outputs a specific policy with maximal performance regardless of the reward signals poisoned by the attack. Then, all attacks that satisfy the budget constraints are ‘optimal’ solutions. As a result, it is meaningless for the attacker to consider all possible algorithms, especially the ones that are not reward-driven and inefficient. A rational learning agent will only consider using an algorithm that ‘efficiently’ learns the reward signals to find a decent policy based on the dataset. Therefore, it is more important for an attack to focus on the effect of the attack against such algorithms. For this purpose, in this work, we assume $\mathcal{L}$ to include all $\delta$-optimal pessimistic learning algorithms in Assumption 4.1 below.
Assumption 4.1.
@@ 4.1 Universal Attack on General Offline Reinforcement Learning as An Optimization Problem
($\delta$-optimal pessimistic learning algorithms assumption) Given an offline RL dataset $\mathcal{D}$ collected from an environment $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R})$ with state-action distribution $\mu$. For all efficient offline RL algorithm $\text{Alg}\in\mathcal{L}$ that might be used by the agent, there exists a small value $\delta>0$ and a class of policies $\Pi_{\mu}$ supported by the distribution $\mu$ such that $J_{\mathcal{R}}(\pi^{\text{Alg}}(\mathcal{D}))\geq\max_{\pi\in\Pi_{\mu}}J_{\mathcal{R}}(\pi)-\delta$. That is, the agent will always use an algorithm that can learn the policy with nearly the highest performance on $\mathcal{M}$ among the supported policies.
Here, we justify that Assumption 4.1 is practical and can represent current efficient offline RL algorithms. It has become common sense that pessimistic learning is the key to efficiency in offline RL Levine et al. (2020). A pessimistic learning algorithm avoids the policies not supported by the dataset to minimize the chance of reward hacking, that is, outputting a policy that appears promising by the dataset but has poor performance. Li et al. (2024) has the same insight in explaining the ‘survival instinct’ of existing offline RL algorithms against attack: a pessimistic algorithm will stick to the policies covered by the dataset so that it won’t output any low-performing and out-of-distribution policies regardless of the reward signals.
By the optimization problem in Eq 1, we can naturally define the efficiency of an attack as follows.
@@
($(V,B,C)$-efficient attack) An attack is $(V,B,C)$-efficient if it satisfies the following conditions:
$\forall\text{Alg}\in\mathcal{L},(\mathcal{D}\oplus\mathbf{\Delta}))\leq V$ (2)
$\|\mathbf{\Delta}\|_{0}\leq B,\|\mathbf{\Delta}\|_{1}\leq C$
Given attack budgets $B$ and $C$, finding the optimal attack for Eq 1 is equivalent to finding the most efficient $(V,B,C)$-efficient attack with the minimal value of $V$. Given an arbitrary attack strategy, the attack budgets $B$ and $C$ can be computed by the attack construction. However, evaluating the learning outcome of the agent under the attack is challenging. First, the influence of a universal attack should be evaluated on all algorithms from $\mathcal{L}$, and it is challenging to find all the algorithms that satisfy the near-optimal efficient learning conditions. Second, evaluating the exact influence of an attack on a specific learning algorithm is challenging in the black-box setting, as the neural network used by the agent is unknown to the attacker. In conclusion, it is challenging to evaluate the exact efficiency of an arbitrary attack; hence, it is also challenging to find the optimal attack. To deal with the challenges and find an attack of high efficiency, we adopt an attack framework where one can analyze the efficiency of the attack instances under the framework.
4.2 Adversarial Reward Engineering Framework
Definition 4.3.
(Adversarial reward engineering framework) In the adversarial reward engineering framework, an instance of attack constructs an adversarial reward function $\widehat{\mathcal{R}}$. The corresponding corruption on the $i^{\text{th}}$ data $(s_{i},a_{i},r_{i})$ from $\mathcal{D}$ satisfies $\Delta_{i}=\widehat{\mathcal{R}}(s_{i},a_{i})-r_{i}$.
In the adversarial reward engineering framework, it is equivalent to say the corrupted dataset is collected from the adversarial environment ^$\widehat{\mathcal{M}}=(\mathcal{S},\mathcal{A},\mathcal{P},\widehat{\mathcal{R}})$ with the same state-action distribution $\mu$ as the original dataset. Therefore, we can apply assumption 4.1 to characterize the behavior of the agent under the attack: it will learn a supported policy that is near-optimal on the adversarial reward function $\widehat{\mathcal{R}}$. Formally, the efficiency of an instance of adversarial reward engineering framework is guaranteed in Lemma 4.4 below.